I need to solve for x in this equation: 3.58x-37.4=.23x+8.32. What does x equal?
Found 2 solutions by midwood_trail, Alan3354:
Answer by midwood_trail(310)

I need to solve for x in this equation: 3.58x -37.4 = .23x + 8.32.
What does x equal?
3.58x -37.4 = .23x + 8.32
Subtract .23x from both sides.
3.58x - .23x -37.4 = 8.32
3.35x - 37.4 = 8.32
Now add 37.4 to both sides.
3.35x = 37.4 + 8.32
3.35x = 45.72
To find x, we now divide both sides of the equation by 3.35.
x = 45.72/3.35
x = 13.65 when rounded off to the nearest tenths.

Answer by Alan3354(69443)

I need to solve for x in this equation: 3.58x-37.4=.23x+8.32. What does x equal?
---------
3.58x-37.4=.23x+8.32
Subtract .23x
3.35x-37.4 = 8.32
Add 37.4
3.35x = 45.72
Divide by 3.35
x = 13.64776...